In the network determine the current flowing through the 4 ohm resistor

5 4 2 w 50230 j2 --j2 50/900
Answer #1

Que. In the Network determine the current flaving through the tr ar resistor 4 2 5 solgu -jz ī 50130 fig.lt) source transformafter redrawing the the fig (2) 4^ (1) 25 10130 Z1 |z2 25 L90 Z1 = fig (3) 5 || ja Shop do 529 121= 10 5²+22 190 Z1 = 5x ja L10 168 196 J2L-45 4r Zz 21 I 6 25 Zz L go = 25x52 1-45 & gó 17 2552 145 10x10 1684198+30 529 98.198 = 100 529 the 4 ohm resinI - 27.6367 -36-636 5.6885 +1 0.722 tant 6.636 2 +27:6367 I= 1627.6367 )+66-6362 115.6885)2+(0.722)? tant {018 00722 2 5-6885